# MaskBit: Embedding-free Image Generation via Bit Tokens

### v1

@@ 3 MaskBit: A New Embedding-free Image Generation Model
Masked “Groups of Bits” Modeling.  During training, the network is supervised with a categorical cross-entropy loss applied to all masked tokens. With increasing number of bits, the categorical cross-entropy gets computed over an exponentially increasing distribution size. Given that bit tokens capture a channel-wise binary quantization, we explore masking “groups of bits”. Specifically, for each bit token $t\in\{-1,1\}^{K}$, we split it into $N$ groups $t_{n}\in\{-1,1\}^{K/N}$, $\forall n\in\{1,\cdots,N\}$, with each group contains $K/N$ consecutive bits. During the masking process, each group of bits can be independently masked. Consequently, a bit token $t$ may be partially masked, allowing the model to leverage unmasked groups to predict the masked bits, easing the training process. During the inference phase, the sampling procedure allows to sample some groups and use their values to guide the remaining samplings. However, this approach increases the number of bit token groups to be sampled, posing a challenge during inference due to the potential for poorly chosen samples. Empirically, we found that using two groups yields the best performance, striking a good balance.
@@ 4 Experimental Results for Image Generation
Experimental Setup.  We evaluate the proposed MaskBit on class-conditional image generation. Specifically, the network generates a total of 50,000 samples for the 1,000 ImageNet [8] classes. All Stage-II generative networks are trained and evaluated on ImageNet at resolutions of $256\times 256$. We report the main metric Fréchet Inception Distance for generation (gFID) [18], along with the side metric Inception Score (IS) [44]. All scores are obtained using the official evaluation suite and reference statistics by ADM [10]. We train all Stage-II networks for 1,080 epochs and report results with classifier-free guidance [21]. More training and inference details can be found in Sec. B.
@@ 4.1 Main Results
Table 1: ImageNet $256\times 256$ generation results. †: Tokenizers trained on OpenImages [28]. ‡: Tokenizers trained on OpenImages, LAION-Aesthetics/-Humans [45]. : Concurrent methods. All results use classifier-free guidance or rejection sampling, but not both. #Params: Generator’s parameters. #Steps: Sampling steps.
@@
LDM-4-G [43] 4096$\times$3 - 3.60 247.7 400M 250
U-ViT [2] 1024$\times$4 - 2.29 263.9 501M 50
DiT-XL/2-G [41] 1024$\times$4 - 2.27 278.2 675M 250
MDT [13] 1024$\times$4 - 1.79 283.0 676M 250
DiMR [35] 1024$\times$4 - 1.70 289.0 505M 250
MDTv2 [14] 1024$\times$4 - 1.58 314.7 676M 250
MAR-H [33] 256$\times$16 - 1.55 303.7 943M 256
@@
Taming-VQGAN [11] 256 1024 5.88 304.8 1.4B 256
MaskGIT [4] 256 1024 4.02 355.6 177M 8
RQ-Transformer [29] 256 16384 3.80 323.7 3.8B 64
ViT-VQGAN [53] 1024 8192 3.04 227.4 1.7B 1024
LlamaGen-3B [47] 576 16384 2.18 263.3 3.1B 576
TiTok-S-128 [57] 128 4096 1.97 281.8 287M 64
VAR [48] 680 4096 1.92 350.2 2.0B 10
MAGVIT-v2 [56] 256 262144 1.78 319.4 307M 64
@@
VQGAN+ as a New Strong Baseline for Image Tokenizers.  By carefully revisiting the VQGAN design in Sec. 2, we develop VQGAN+, a solid new baseline for image tokenizers that achieves an rFID of 1.66, marking a significant improvement of 6.28 rFID over the original Taming-VQGAN model [11]. Using VQGAN+ in the generative framework of MaskGIT [4] results in a gFID of 2.12, outperforming several recent advanced VQGAN-based methods, such as ViT-VQGAN [53] (by 0.92 gFID), RQ-Transformer [29] (by 1.68 gFID), and even the original MaskGIT [4] (by 1.90 gFID). This result shows that the improvements we propose to the Stage-I model, indeed transfer to the generation model. Furthermore, our result surpasses several recent diffusion-based generative models that utilize the heavily optimized VAE [27] by Stable Diffusion [1], including LDM [43], U-ViT [2], and DiT [41]. We believe these results with VQGAN+ establish a new solid baseline for the community.
Bit Tokens as a Strong Representation for Both Reconstruction and Generation.  As discussed in previous sections, bit tokens encapsulate a structured semantic representation that retains essential information for image generation, forming the foundation of MaskBit. Consequently, MaskBit delivers superior performance, achieving a gFID of 1.65 with 12 bits and 1.62 with 14 bits, surpassing prior works. Notably, MaskBit outperforms the recent state-of-the-art MAGVIT-v2 [56] while utilizing an implicit codebook that is 16×16$\times$16 × to 64×64$\times$64 × smaller. When the number of sampling steps is increased from 64 to 256—matching the steps used by autoregressive models—MaskBit sets a new state-of-the-art with a gFID of 1.52, outperforming concurrent works [14, 33, 35, 47, 48, 57], while using a compact generator model of just 305M parameters. Remarkably, MaskBit, which operates with discrete tokens, surpasses methods relying on continuous tokens, such as MDTv2 [14] and MAR-H [33].
Table 2: Comparison using similar generator model capacity ($\sim$ 300M) on ImageNet $256\times 256$. †: Tokenizers trained on OpenImages [28]. : Concurrent methods. All results use classifier-free guidance. #Steps: Sampling steps.
@@
Comparison under Similar Generator Parameters.  In Tab. 2, we present a performance comparison across models of similar sizes. Previous works have demonstrated that increasing model size generally leads to superior performance [33, 47, 48]. Therefore, to provide a more comprehensive comparison, we only compare models using a similar number of parameters, but vary codebook size and sampling steps. As shown in the table, MaskBit not only achieves state-of-the-art results overall but also outperforms all other models under similar conditions. It is worth noting that all models in the comparison use a larger Stage-I (reconstruction) tokenizer than ours.
Qualitative Results.  We present visualizations of the generated $256\times 256$ images using the proposed MaskBit in Fig. 5. The results show the diversity of classes represented in ImageNet. MaskBit is able to generate high-fidelity outputs. We note that these generated results inherit the dataset bias of ImageNet. As a classification dataset, the images are centered around objects showing the respective class. Thus, this dataset bias influences the kind of images that can be generated. Methods used in production environments such as Stable Diffusion [1] therefore train on more general large-scale datasets such as LAION-5B [45] to remove such dataset bias.
@@ 4.2 Ablation Study
Unless stated otherwise, we use fewer training iterations for the ablation studies due to the limited resources and computational costs.
Embedding-free Stages Bring Improvements.  In Tab. 3(a), we investigate embedding-free designs in both Stage-I (tokenizer) and Stage-II (generator). To ensure a fair comparison, all models use the same number of tokens, codebook size, and model capacity. In the first row, we employ VQGAN+ with a traditional Stage-II model, as in MaskGIT [4], achieving a gFID of 2.12. Using bit tokens in Stage-I improves performance to a gFID of 1.95, highlighting the effectiveness of structured semantic representation of bit tokens. Further enabling our Stage-II model MaskBit (i.e., both stages now are embedding-free) achieves the best gFID of 1.82, demonstrating the advantages of embedding-free image generation from bit tokens. For completeness, we also explore VQGAN+ with an embedding-free Stage-II model, yielding a gFID of 2.83. This variant performs worse due to the lack of the semantic structure in the tokens.
Two Groups of Bits Lead to Better Performance.   Tab. 3(b) presents the performance of MaskBit using a different numbers of groups. As shown in the table, when using two groups, each consisting of six consecutive bits, MaskBit achieves the best performance. This configuration outperforms a single group because it allows the model to leverage information from the unmasked groups. However, when using four groups, the increased difficulty in sampling during inference – due to the need for four times more sampling – detracts from performance. Still, all group configurations outperform our strong baseline using VQGAN+ as well as modern approaches like ViT-VQGAN [53] and DiT-XL [41], showing the potential of grouping. We also conduct another experiment, where we apply 2 groups to VQGAN+. The performance drops form 2.12 to 4.4 gFID, due to the missing semantic structure in the tokens.
@@
14 Bits are Enough for ImageNet.  In Tab. 4(a), we train the Stage-II models with our final training schedule. In each setting, MaskBit leverages a different number of bits. Empirically, we find that using 14 bits works the best on ImageNet. We note that all settings outperform prior methods (e.g.[41, 56]) in gFID and the performance differences are marginal across different number of bits. While using 14 bits is sufficient for ImageNet, more general and large-scale datasets such as LAION-5B [45] might benefit from using larger values of $K$ (i.e., larger implicit vocabulary sizes).
MaskBit Scales with More Sampling Steps.  MaskBit follows the non-autoregressive sampling paradigm [4, 55], enabling flexibility in the number of sampling steps during inference (up to 256 steps in our ImageNet 256$\times$256 experiments). Unlike autoregressive models [11, 47], this approach allows for fewer forward passes through the Stage-II generative model, reducing computational cost and inference time. However, increasing MaskBit’s sampling steps to match those of autoregressive models can also improve performance. As shown in Tab. 4(b), even with 64 steps, MaskBit surpasses prior work and achieves an excellent balance between compute efficiency and performance. Scaling further to 256 steps sets a new state-of-the-art gFID of 1.52.
@@ 5 Related Work
Image Tokenization. Images represented by raw pixels are highly redundant and commonly compressed to a latent space with autoencoders [20, 51], since the early days of deep learning. Nowadays, image tokenization still plays a crucial role in mapping pixels into discrete [40] or continuous [27, 19] representation suitable for generative modeling. We focus on the stream of vector quantization (VQ) tokenizers, which proves successful in state-of-the-art transformer generation frameworks [11, 4, 54, 56, 31, 32, 29, 30]. Starting from the seminal work VQ-VAE [40], several key advancements have been proposed, such as the perceptual loss [58] and discriminator loss [26] from [11] or better quantization methods [60, 29, 59, 39, 56], which significantly improve the reconstruction quality. Nonetheless, modern transformer generation models [4, 55, 56] utilize an even stronger VQ model, with details not fully discussed in the literature, resulting in a performance gap to the wider research community. In this work, we conducted a systematic study to modernize a widely-used VQGAN model, aiming at demystifying the process to obtain a strong tokenizer.
Image Generation. Although various types of image generation models exist, mainstream state-of-the-art methods are usually either diffusion-based [10, 7, 43, 41, 34, 23], auto-regressive-based [6, 11, 53, 12], or masked-transformer-based [4, 30, 56, 31, 32]. Among them, the masked-transformer-based methods exhibit competitive performance, while bringing a substantial speed-up, thanks to a much fewer sampling steps. This research line begins with MaskGIT [4], which generates images from a masked sequence in non-autoregressive manner where at each step multiple token can be predicted [9]. Follow-up work has successfully extend the idea to the video generation domain [55, 16], large-scale text-to-image generation [5], along with competitive performance to diffusion models [56] for image generation. Nonetheless, these works only utilize the tokens from the tokenizer to learn new embeddings per token in the transformer network, while we aim at an embedding-free generator framework, making better use of the learned semantic structure from the tokenizer.
@@ 6 Conclusion
  • Sec. C provides more discussion on bit tokens.
  • Sec. D provides more reconstructed images for the bit flipping experiments.
  • Sec. E discusses the method’s limitations.
  • Sec. F discusses the broader impacts.
@@ Appendix A Datasets
ImageNet: ImageNet [8] is one of the most popular benchmarks in computer vision. It has been used to benchmark image classification, class-conditional image generation, and more.

### v2

@@ 3 MaskBit: A New Embedding-free Image Generation Model
Masked “Groups of Bits” Modeling.  During training, the network is supervised with a categorical cross-entropy loss applied to all masked tokens. With increasing number of bits, the categorical cross-entropy gets computed over an exponentially increasing distribution size. Given that bit tokens capture a channel-wise binary quantization, we explore masking “groups of bits”. Specifically, for each bit token $t\in\{-1,1\}^{K}$, we split it into $N$ groups $t_{n}\in\{-1,1\}^{K/N}$, $\forall n\in\{1,\cdots,N\}$, with each group contains $K/N$ consecutive bits. During the masking process, each group of bits can be independently masked. Consequently, a bit token $t$ may be partially masked, allowing the model to leverage unmasked groups to predict the masked bits, easing the training process. An illustration of the grouping and masking procedure can be found in Appendix C. During the inference phase, the sampling procedure allows to sample some groups and use their values to guide the remaining samplings. However, this approach increases the number of bit token groups to be sampled, posing a challenge during inference due to the potential for poorly chosen samples. Empirically, we found that using two groups yields the best performance, striking a good balance.
@@ 4 Experimental Results for Image Generation
Experimental Setup.  We evaluate the proposed MaskBit on class-conditional image generation. Specifically, the network generates a total of 50,000 samples for the 1,000 ImageNet (Deng et al., 2009) classes. All Stage-II generative networks are trained and evaluated on ImageNet at resolutions of $256\times 256$. We report the main metric Fréchet Inception Distance for generation (gFID) (Heusel et al., 2017), along with the side metric Inception Score (IS) (Salimans et al., 2016). All scores are obtained using the official evaluation suite and reference statistics by ADM (Dhariwal & Nichol, 2021). We train all Stage-II networks for 1,080 epochs and report results with classifier-free guidance (Ho & Salimans, 2022). More training and inference details can be found in Sec. B.
@@ 4.1 Main Results
Table 1: ImageNet $256\times 256$ generation results. †: Tokenizers trained on OpenImages (Kuznetsova et al., 2020). ‡: Tokenizers trained on OpenImages, LAION-Aesthetics/-Humans (Schuhmann et al., 2022). : Concurrent methods. All results use classifier-free guidance or rejection sampling, but not both. #Params: Generator’s parameters. #Steps: Sampling steps.
@@
LDM-4-G (Rombach et al., 2022) 4096$\times$3 - 3.60 247.7 400M 250
GIVT-L-A (Tschannen et al., 2024) 256$\times$32 - 2.59 - 1.67B 256
U-ViT (Bao et al., 2023) 1024$\times$4 - 2.29 263.9 501M 50
DiT-XL/2-G (Peebles & Xie, 2023) 1024$\times$4 - 2.27 278.2 675M 250
MDT (Gao et al., 2023) 1024$\times$4 - 1.79 283.0 676M 250
DiMR (Liu et al., 2024) 1024$\times$4 - 1.70 289.0 505M 250
MDTv2 (Gao et al., 2024) 1024$\times$4 - 1.58 314.7 676M 250
MAR-H (Li et al., 2024) 256$\times$16 - 1.55 303.7 943M 256
@@
Taming-VQGAN (Esser et al., 2021) 256 1024 5.88 304.8 1.4B 256
MaskGIT (Chang et al., 2022) 256 1024 4.02 355.6 177M 8
RQ-Transformer (Lee et al., 2022a) 256 16384 3.80 323.7 3.8B 64
ViT-VQGAN (Yu et al., 2022a) 1024 8192 3.04 227.4 1.7B 1024
LlamaGen-3B (Sun et al., 2024) 576 16384 2.18 263.3 3.1B 576
TiTok-S-128 (Yu et al., 2024c) 128 4096 1.97 281.8 287M 64
VAR (Tian et al., 2024) 680 4096 1.92 350.2 2.0B 10
MAGVIT-v2 (Yu et al., 2024a) 256 262144 1.78 319.4 307M 64
@@
VQGAN+ as a New Strong Baseline for Image Tokenizers.  By carefully revisiting the VQGAN design in Sec. 2, we develop VQGAN+, a solid new baseline for image tokenizers that achieves an rFID of 1.66, marking a significant improvement of 6.28 rFID over the original Taming-VQGAN model (Esser et al., 2021). Using VQGAN+ in the generative framework of MaskGIT (Chang et al., 2022) results in a gFID of 2.12, outperforming several recent advanced VQGAN-based methods, such as ViT-VQGAN (Yu et al., 2022a) (by 0.92 gFID), RQ-Transformer (Lee et al., 2022a) (by 1.68 gFID), and even the original MaskGIT (Chang et al., 2022) (by 1.90 gFID). This result shows that the improvements we propose to the Stage-I model, indeed transfer to the generation model. Furthermore, our result surpasses several recent diffusion-based generative models that utilize the heavily optimized VAE (Kingma & Welling, 2014) by Stable Diffusion (AI, 2022), including LDM (Rombach et al., 2022), U-ViT (Bao et al., 2023), and DiT (Peebles & Xie, 2023). We believe these results with VQGAN+ establish a new solid baseline for the community.
Bit Tokens as a Strong Representation for Both Reconstruction and Generation.  As discussed in previous sections, bit tokens encapsulate a structured semantic representation that retains essential information for image generation, forming the foundation of MaskBit. Consequently, MaskBit delivers superior performance, achieving a gFID of 1.65 with 12 bits and 1.62 with 14 bits, surpassing prior works. Notably, MaskBit outperforms the recent state-of-the-art MAGVIT-v2 (Yu et al., 2024a) while utilizing an implicit codebook that is 16×16$\times$16 × to 64×64$\times$64 × smaller. When the number of sampling steps is increased from 64 to 256—matching the steps used by autoregressive models—MaskBit sets a new state-of-the-art with a gFID of 1.52, outperforming concurrent works (Gao et al., 2024; Li et al., 2024; Liu et al., 2024; Sun et al., 2024; Tian et al., 2024; Yu et al., 2024c), while using a compact generator model of just 305M parameters. Remarkably, MaskBit, which operates with discrete tokens, surpasses methods relying on continuous tokens, such as MDTv2 (Gao et al., 2024) and MAR-H (Li et al., 2024).
Table 2: Comparison using similar generator model capacity ($\sim$ 300M) on ImageNet $256\times 256$. †: Tokenizers trained on OpenImages (Kuznetsova et al., 2020). : Concurrent methods. All results use classifier-free guidance. #Steps: Sampling steps.
@@
Comparison under Similar Generator Parameters.  In Tab. 2, we present a performance comparison across models of similar sizes. Previous works have demonstrated that increasing model size generally leads to superior performance (Li et al., 2024; Sun et al., 2024; Tian et al., 2024). Therefore, to provide a more comprehensive comparison, we only compare models using a similar number of parameters, but vary codebook size and sampling steps. As shown in the table, MaskBit not only achieves state-of-the-art results overall but also outperforms all other models under similar conditions. It is worth noting that all models in the comparison use a larger Stage-I (reconstruction) tokenizer than ours.
Qualitative Results.  We present visualizations of the generated $256\times 256$ images using the proposed MaskBit in Fig. 5. The results show the diversity of classes represented in ImageNet. MaskBit is able to generate high-fidelity outputs. We note that these generated results inherit the dataset bias of ImageNet. As a classification dataset, the images are centered around objects showing the respective class. Thus, this dataset bias influences the kind of images that can be generated. Methods used in production environments such as Stable Diffusion (AI, 2022) therefore train on more general large-scale datasets such as LAION-5B (Schuhmann et al., 2022) to remove such dataset bias.
@@ 4.2 Ablation Study
Unless stated otherwise, we use fewer training iterations for the ablation studies due to limited resources and computational costs.
Embedding-free Stages Bring Improvements.  In Tab. 3(a), we investigate embedding-free designs in both Stage-I (tokenizer) and Stage-II (generator). To ensure a fair comparison, all models use the same number of tokens, codebook size, and model capacity. In the first row, we employ VQGAN+ with a traditional Stage-II model, as in MaskGIT (Chang et al., 2022), achieving a gFID of 2.12. Using bit tokens in Stage-I improves performance to a gFID of 1.95, highlighting the effectiveness of structured semantic representation of bit tokens. Further enabling our Stage-II model MaskBit (i.e., both stages now are embedding-free) achieves the best gFID of 1.82, demonstrating the advantages of embedding-free image generation from bit tokens. For completeness, we also explore VQGAN+ with an embedding-free Stage-II model, yielding a gFID of 2.83. This variant performs worse due to the lack of the semantic structure in the tokens.
Two Groups of Bits Lead to Better Performance.   Tab. 3(b) presents the performance of MaskBit using a different numbers of groups. As shown in the table, when using two groups, each consisting of six consecutive bits, MaskBit achieves the best performance. This configuration outperforms a single group because it allows the model to leverage information from the unmasked groups. However, when using four groups, the increased difficulty in sampling during inference – due to the need for four times more sampling – detracts from performance. Still, all group configurations outperform our strong baseline using VQGAN+ as well as modern approaches like ViT-VQGAN (Yu et al., 2022a) and DiT-XL (Peebles & Xie, 2023), showing the potential of grouping. We also conduct another experiment, where we apply 2 groups to VQGAN+. The performance drops form 2.12 to 4.4 gFID, due to the missing semantic structure in the tokens.
@@
MaskBit with 14 Bits Yields Better Performance.  In Tab. 4(a), we train the Stage-II models with our final training schedule. In each setting, MaskBit leverages a different number of bits. Empirically, we find that using 14 bits works best on ImageNet. We note that all settings outperform prior methods (e.g.Peebles & Xie (2023); Yu et al. (2024a)) in gFID and the performance differences are marginal across different number of bits. While using 14 bits is sufficient for ImageNet, more general and large-scale datasets such as LAION-5B (Schuhmann et al., 2022) might benefit from using more bits (i.e., larger implicit vocabulary sizes).
MaskBit Scales with More Sampling Steps.  MaskBit follows the non-autoregressive sampling paradigm (Chang et al., 2022; Yu et al., 2023), enabling flexibility in the number of sampling steps during inference (up to 256 steps in our ImageNet 256$\times$256 experiments). Unlike autoregressive models (Esser et al., 2021; Sun et al., 2024), this approach allows for fewer forward passes through the Stage-II generative model, reducing computational cost and inference time. However, increasing MaskBit’s sampling steps to match those of autoregressive models can also improve performance. As shown in Tab. 4(b), even with 64 steps, MaskBit surpasses prior work and achieves an excellent balance between compute efficiency and performance. Scaling further to 256 steps sets a new state-of-the-art gFID of 1.52.
@@ 5 Related Work
Image Tokenization. Images represented by raw pixels are highly redundant and commonly compressed to a latent space with autoencoders (Hinton & Salakhutdinov, 2006; Vincent et al., 2008), since the early days of deep learning. Nowadays, image tokenization still plays a crucial role in mapping pixels into discrete (Oord et al., 2017) or continuous (Kingma & Welling, 2014; Higgins et al., 2017) representation suitable for generative modeling. We focus on the stream of vector quantization (VQ) tokenizers, which proves successful in state-of-the-art transformer generation frameworks (Esser et al., 2021; Chang et al., 2022; Yu et al., 2022b; 2024a; Lezama et al., 2022; 2023; Lee et al., 2022a; b). Starting from the seminal work VQ-VAE (Oord et al., 2017), several key advancements have been proposed, such as the perceptual loss (Zhang et al., 2018) and discriminator loss (Karras et al., 2019) from (Esser et al., 2021) or better quantization methods (Zheng et al., 2022; Lee et al., 2022a; Zheng & Vedaldi, 2023; Mentzer et al., 2024; Yu et al., 2024a), which significantly improve the reconstruction quality. Nonetheless, modern transformer generation models (Chang et al., 2022; Yu et al., 2023; 2024a) utilize an even stronger VQ model, with details not fully discussed in the literature, resulting in a performance gap to the wider research community. In this work, we conducted a systematic study to modernize the VQGAN model, aiming at demystifying the process to obtain a strong tokenizer.
Image Generation. Although various types of image generation models exist, mainstream state-of-the-art methods are usually either diffusion-based (Dhariwal & Nichol, 2021; Chen et al., 2023; Rombach et al., 2022; Peebles & Xie, 2023; Li et al., 2023; Hoogeboom et al., 2023), auto-regressive-based (Chen et al., 2020; Esser et al., 2021; Yu et al., 2022a; Gafni et al., 2022; Tschannen et al., 2024; Yu et al., 2024b), or masked-transformer-based (Chang et al., 2022; Lee et al., 2022b; Yu et al., 2024a; Lezama et al., 2022; 2023). Among them, the masked-transformer-based methods exhibit competitive performance, while bringing a substantial speed-up, thanks to a much fewer sampling steps. This research line begins with MaskGIT (Chang et al., 2022), which generates images from a masked sequence in non-autoregressive manner where at each step multiple token can be predicted (Devlin et al., 2018). Follow-up work has successfully extend the idea to the video generation domain (Yu et al., 2023; Gupta et al., 2023), large-scale text-to-image generation (Chang et al., 2023), along with competitive performance to diffusion models (Yu et al., 2024a) for image generation. Still, these works only utilize the tokens from the tokenizer to learn new embeddings in the transformer network, while we aim at an embedding-free generator framework, making better use of the learned semantic structure from the tokenizer.
@@ 6 Conclusion
  • Sec. C provides a visual explanation on masking groups of bit tokens.
  • Sec. D provides more discussion on bit tokens.
  • Sec. E provides more reconstructed images for the bit flipping experiments on ImageNet and COCO.
  • Sec. F provides zero-shot reconstruction results on COCO.
  • Sec. G provides a nearest-neighbor analysis on ImagetNet.
  • Sec. H discusses the method’s limitations.
  • Sec. I discusses the broader impacts.
@@ Appendix A Datasets
ImageNet: ImageNet (Deng et al., 2009) is one of the most popular benchmarks in computer vision. It has been used to benchmark image classification, class-conditional image generation, and more.